# A uniform kernel trick for high and infinite-dimensional two-sample problems

### v2

@@ 4 Empirical results
In the same spirit as Borgwardt et al. (2012) or (Fukumizu et al., 2007, Section 8.1), we emphasize the interest of a new homogeneity test (based on kernel distances), suitable for high-dimensional data and not suffering from the degradation of classical two-sample tests when the dimension increases. Additionally, we show the advantages of avoiding the choice of the parameters in kernel distances, via our SKD proposal. The general idea is to check the SKD methodology as an attempt to robustify the test statistic against bad choices of the kernel or its parameter(s). In this empirical study we compare the following methods:
@@
GKD: the kernel distance-based test of Fukumizu et al. (2007) with a data-driven choice of $\lambda$ in (2) as the median distance between points in the aggregate sample.
GKDSplit: test based on a Gaussian kernel distance where the estimation of the parameters is done by a splitting method; see Fukumizu et al. (2009). The sample is divided into training and test subsamples to avoid data influence on the parameter selection.
GKDSplitOpt: test based on a kernel distance where the parameter estimation is detailed in (Balakrishnan et al., 2012, Section 3). The data is divided into training and test sets. The target parameters are the coefficients of a convex combination of a finite family of kernels. The weights of the combination are selected to maximize the ratio between the empirical bias-corrected kernel distance and the standard deviation of the associated asymptotic distribution of the centered, normalized empirical distance.
ET: the energy test, a popular choice in this type of problems; see Rizzo and Székely (2017), Rizzo and Szekely (2022).
In our view, the “data-splitting” proposals are based on natural ideas that deserve attention. Still, there are some open issues to clarify, especially regarding the optimal splitting of the sample and the asymptotic behaviour of the resulting data-driven tests. We hope that this study could encourage further research along these lines. As for the energy test ET (Rizzo and Székely (2017)), we have included it in the study because it is based in a successful statistical methodology, ultimately grounded on the underlying “distance covariance” association measure; in fact this method has become quite popular in high-dimensional two sample problems, via the energy package in R. Following a suggestion from one of the reviewers, we have considered as well a variant of this method, which is based on a different distance between the sample points. The standard statistic in ET is calculated in terms of the Euclidean distance, which can be seen, by a duality reasoning explained in Rizzo and Székely (2017), as a distance associated with the Brownian covariance. This suggests the possibility of broadening the choice of the distance to the whole range of the fractional Brownian motion with Hurst parameter $H\in(0,1)$; recall that $H={1}/{2}$ for the standard Brownian motion. In our case, the choice $H={3}/{4}$ led to results almost identical to those of $H={1}/{2}$. Perhaps if heavy-tailed distributions were involved, this new alternative could make a real difference.
@@
We include the models in Fukumizu et al. (2007), based on Gaussian distributions in high dimension with different means and diagonal covariance matrices. In addition, we also consider a new scenario with functional data corresponding to trajectories of Gaussian processes in $\operatorname{L}^{2}([0,1])$. We note that all the considered tests can be applied in the functional setting: GKD, SKD, GKDSplit and GKDSplitOpt are based on the aggregated matrix $\left(k_{\lambda}\left(Z_{i},Z_{j}\right)\right)_{i,j=1}^{n+m}$, where $Z_{l}=X_{l}$ for $l\in\{1,\ldots,n\}$ and $Z_{n+l}=Y_{l}$ for $l\in\{1,\ldots,m\}$. ET uses cross-distances between the data in the sample space. More specifically, our simulation experiments are grouped in three blocks, respectively corresponding to different versions of homoscedasticity (Experiment 1) and heteroscedasticity (Experiment 2), plus a functional real data example.
@@
This regularization is common in harmonic analysis to approximate the Dirac delta in spaces of distributions via smooth functions, called mollifiers. It could be seen as an ad-hoc correction to improve the approximation of the maximum of the estimated kernel distance to the corresponding “true” population maximum. The smoothing process can be eliminated when sample sizes are sufficiently large. As shown in the literature, a data-driven choice of $\lambda$ seems to have a good practical behaviour. Specifically, in Fukumizu et al. (2007) (and in subsequent works), the value of $\lambda$ is the median distance between points in the aggregate sample. A theoretical consequence of this choice is that the asymptotic theory, derived in Fukumizu et al. (2007) under the assumption that $\lambda$ is fixed, does not longer apply to the data-driven case. Still, we include in our experiments, for comparison purposes, this data-driven choice as it is a common practice in the earlier literature. Since, to the best of our knowledge, the asymptotic distribution of the data-driven statistic is not known, we use a permutation test based on this statistic to obtain rejection regions rather than the other methods (Pearson curves, Gamma curves and bootstrap for U-statistics) explained in Fukumizu et al. (2007).
@@
It is worth mentioning that in both tests SKD and ET a permutation procedure has been used to approximate the corresponding distributions. This is the methodology used in the energy R-package for the ET test and we have followed here the the same strategy for the SKD test. Let us recall that our theoretical results provide the asymptotic distribution of this test, as well as its consistency (see Theorems 2 and 3). However, the estimation of the quantiles of the limit distribution in (13) is far from trivial. As an additional complication, standard bootstrap approximation fails, as a consequence of the results in Fang and Santos (2019). This is why the permutation method appears as a natural choice.
@@
Let us recall that the idea behind the SKD test is to dodge the parameter selection problem by considering “the whole parameter space”. Ideally, for the Gaussian kernel, an interval of the form $(0,\infty)$ could be considered in the SKD. As mentioned before, the extremes (00 and $\infty$) are not useful since the distance tends to zero when the parameter approaches to these values. Therefore, we use a parameter space of the form $\Lambda=[a,b]$, with $0<a<b<\infty$. It is important to note that here the values $a$ and $b$ cannot be properly considered as tuning parameters, since the test is not particularly sensitive to their choice, provided that the interval is large enough. As we have experimentally verified, $\Lambda=\left[10^{-4},0.1\right]$ is adequate to carry out the test. In practice, we employ a grid of $11$ points logarithmically separated between $10^{-4}$ and $0.1$. The goal is to approximate the value of the supremum by the maximum over finite subsets. The simulation outputs below are based on averages over $200$ replications. The permutation tests for GKD and SKD correspond to $B=5000$ permutations. As for ET, we use the function eqdist.etest of the R-package Rizzo and Szekely (2022). Sample sizes are $n=m=250$ in all experiments. The effect of increasing the dimension $d$ is checked in the rank $d\in\{205,405,603,803,1003,1203,1401,$ 1601,1801,2001}1601,1801,2001\}1601 , 1801 , 2001 }. In all cases, the significance level of the test is set at $\alpha=0.05$.
@@
We consider daily values of maximum temperatures registered at Barcelona airport (El Prat) from years 1944 to 2019. The data set consists of 76 vectors of dimension 365, each of which corresponds to a year in that time period. The daily observations have been treated as discretization points to include the problem within the framework of functional data, every year providing a function in the sample. Those observations corresponding to the 29th of February in leap years are omitted and missing observations are interpolated. These data are available at https://www.ncei.noaa.gov, the web page of the National Centers for Environmental Information. Our purpose is to test the null hypothesis that the sample of temperatures from 1944 to 1981 comes from the same (functional) distribution to that of the period 1982-2019. The rejection of this null hypothesis could be interpreted as a hint of possible warming in the area. Indeed, we observe that, in absence of any significant climate change, one would expect that both samples are made of independent trajectories from the same underlying process.
@@
In the light of the results, we can conclude that, globally, the supremum kernel distance test (SKD) performs similarly to the GKD tests in the homoscedastic case, though the ET test appears to be the winner in this situation. In the heteroscedastic case, SKD obtains almost the best power results. A more complete study (including the derivation of the asymptotic distribution for the case of a data-driven selection of $\lambda$, the use of Pearson curves and/or modified bootstrap schemes, …) might be worthwhile in the future. On the other hand, according to Fukumizu et al. (2013), energy tests can be expressed in terms of kernel distances. This idea might deserve further attention as well, in order to incorporate these “equivalent” kernels to the SKD paradigm. In any case, it is clear that the present study does not allow us to conclude any obvious superiority (or inferiority) of none of the considered methods. In fact, the aim of our limited empirical study is to show that the SKD method can be implemented and it is competitive. This goal has been hopefully achieved. More definitive conclusions should be reached via subsequent empirical experiments and, especially, with the use of these tests by practitioners in the coming years. Software to run the SKD-based test will soon be available as an R-package called SKD2.
@@ 5 Final remarks
The whole paper relies on the strategy of using kernel-based methods by combining, via supremum, different kernels. We consider the classical two-sample problem focusing on functional and high-dimensional data, where the demand for applicable methods seems more obvious. Despite the large amount of relevant literature on this topic, we believe that there is still room for improvement, as those provided here, in the line of obtaining more general results with a different technology of proofs. In particular, the use of differentiation techniques plus empirical processes methods allows us to address the asymptotic behaviour of the test statistics under the null and alternative hypothesis, including the case of unbalanced samples; see Theorems 2 and 3. A key result in the proofs is Theorem 1 that establishes, under some conditions covering infinite-dimensional settings, the universal Donsker property for a union of unit balls in RKHS. This theorem extends previous similar results for finite-dimensional situations and could be potentially useful in other statistical procedures within the RKHS framework. The approach established in this paper can also be potentially useful to analyze the asymptotic behavior of data-driven estimators of the kernel parameters. However, this interesting problem is beyond the scope of this paper. Theorems 2 and 3 are meaningful from a conceptual point of view, even if the limit distributions are not particularly simple. The mere existence of such limit (non-degenerate) distributions is a primary guarantee that kernel-based statistics can be used to derive procedures achieving, at least asymptotically, a prescribed significance level under the null and providing consistency under the alternative. We deal with the complicated structure of the limit distribution under the null by using permutation tests. This is also the case for other popular methodologies such as the energy test, Rizzo and Székely (2017). Different approximation techniques are also conceivable, including truncation in the limit expression in Theorems 2 plus estimation of the involved parameters. While this paper was under review, other related and interesting contributions have appeared; see Albert et al. (2023) and Biggs et al. (2023).
@@ 6 Technical details
We need two auxiliar lemmata to prove Theorem 1. The first result corresponds to (Marcus, 1985, Theorem 1.1).
@@
We also require Aronszajn’s inclusion theorem; see (Aronszajn, 1950, Theorem I).
@@
The first part can be seen as a consequence of Lemma 1. First, by (Berlinet and Thomas-Agnan, 2011, Theorem 17), the functions in $\mathcal{H}_{k}$ are continuous. In particular, using (Berlinet and Thomas-Agnan, 2011, Corollary 3), we conclude that $\mathcal{H}_{k}$ is a separable Hilbert space. On the other hand, for $x\in\mathcal{X}$, by the reproducing property (6) of $k$ (twice) and Cauchy–Schwarz inequality, we have that
@@
Therefore, as $k$ is bounded on the diagonal, convergence in the RKHS norm entails uniform convergence. Further, from (19) we also see that the functions in $\mathcal{H}_{k}$ are bounded and hence $\mathcal{H}_{k}\subset\mathcal{C}_{b}(\mathcal{X})$. Now, we can apply Lemma 1 to $H=\mathcal{H}_{k}$ and $T=I$, the inclusion map given by $I(f)=f$. According to (19), this linear transformation is continuous. As $B_{H}=\mathcal{F}_{k}$, by Lemma 1, we thus conclude that $\mathcal{F}_{k}=T\left(\mathcal{F}_{k}\right)$ is universal Donsker. The second part is a by-product of the first one together with Aronszajn’s inclusion theorem. According to Lemma 2, we have that $\|f\|_{\mathcal{H}_{k}}\leq\sqrt{c}\,\|f\|_{\mathcal{H}_{k,\lambda}}$, for all $f\in\mathcal{H}_{k,\lambda}$ and for all $\lambda\in\Lambda$. Therefore, $\mathcal{F}_{k,\Lambda}\subset\sqrt{c}\,\mathcal{F}_{k}$. Finally, from the first part of the theorem, the set $\sqrt{c}\,\mathcal{F}_{k}$ is universal Donsker as it is the unit ball of the RKHS generated by the kernel $ck$. Therefore, $\mathcal{F}_{k,\Lambda}$ is also universal Donsker (see (van der Vaart and Wellner, 2023, Theorem 2.10.1)) and the proof is complete. ∎
@@
Proof.
To show part ((a)), we note that, from Theorem 1, $\mathcal{F}_{k,\lambda}$ is a $\operatorname{P}$-Donsker class and hence $\operatorname{P}$-pre-Gaussian. Hence, by (Giné and Nickl, 2016, Theorem 3.7.28), for almost all $\omega$, the function $f\mapsto\mathbb{G}_{\operatorname{P}}(\omega)f$ ($f\in\mathcal{F}_{k,\lambda}$) is prelinear and can be uniquely extended to a linear map on $\operatorname{span}\left(\mathcal{F}_{k,\lambda}\right)=\mathcal{H}_{k,\lambda}$. Moreover, this extension is bounded and uniformly $d_{\operatorname{P}}$-continuous in $\mathcal{H}_{k,\lambda}$, where $d_{\operatorname{P}}$ is the intrinsic $\operatorname{L}^{2}(\operatorname{P})$ metric of the process. Finally, we observe that, thanks to (19),
@@
To prove part ((b)) we first note that, by ((a)), $\mathbb{G}_{\operatorname{P}}$ is a Gaussian process in the Hilbert space $\mathcal{H}_{k,\lambda}^{\ast}$. The covariance operator $\mathcal{K}_{\mathbb{G}_{\operatorname{P}}}$ of $\mathbb{G}_{\operatorname{P}}$ is self-adjoint and compact. By the Fernique’s theorem (Bogachev, 1998, p. 74), $\mathbb{G}_{\operatorname{P}}$ is Bochner square-integrable, $\mathcal{K}_{\mathbb{G}_{\operatorname{P}}}$ is a trace-class operator and
@@
where $\nu_{\mathbb{G}_{\operatorname{P}}}$ is the measure induced by the process $\mathbb{G}_{\operatorname{P}}$ in $\mathcal{H}_{k,\lambda}^{\ast}$. The proof of (23) can be found in (Bogachev, 1998, p. 48).
@@
From (24), we have that $Z_{j,\lambda}=\left\langle\mathbb{G}_{\operatorname{P}},\varphi_{j,\lambda}\right\rangle_{\mathcal{H}_{k,\lambda}^{\ast}}\sim\mathcal{N}\left(0,\beta_{j,\lambda}\right)$ ($j\in\mathbb{N}$) are jointly Gaussian and independent. To finish this proof of (20), by (Ledoux and Talagrand, 2013, Theorem 6.1), it is enough to show absolute mean convergence, which is a necessary and sufficient condition. First, by orthogonality, we observe that for every $J\subset\mathbb{N}$ finite, we have that
@@
The proof of part ((a)) in Lemma 3 essentially follows from Theorem 1. However, part ((b)), where the series representation is obtained, must be discussed. Equation (20) shows the convergence of a series of functional random variables. This result looks like a standard Karhunen-Loève theorem, but some remarks should be done. The convergence of this series is on the dual space $\mathcal{H}_{k,\lambda}^{\ast}$, while Karhunen-Loève decomposition is stated classically on $\operatorname{L}^{2}$-type spaces. In fact, our decomposition in (20) can be seen as a particular case of the results in Bay and Croix (2017). In (Giné and Nickl, 2016, Theorem 2.6.10) a similar decomposition is shown where the coordinates are deterministic while the basis is random, which is not useful for our purposes.
@@
Note that $d_{k,\Lambda}$ is the metric induced by the supremum norm in $\ell^{\infty}\left(\mathcal{F}_{k,\Lambda}\right)$, hence $d_{k,\Lambda}$ is a continuous functional. From (26) and by the continuous mapping theorem (see, for instance (van der Vaart and Wellner, 2023, Theorem 1.9.5)), we obtain that
@@
The next goal is to prove Corollary 1 as preparation for the proof of Theorem 3. We need a differentiability result for the supremum similar to those obtained in Cárcamo et al. (2020). Given a kernel $k$, we consider the mapping
@@
where $\mathcal{F}_{k,\Lambda}$ is the union of balls in (11). These differentiability results might have independent interest as it can be applied in other contexts by means of the (extended) functional Delta method; see the examples in Cárcamo et al. (2020). The next corollary shows that $\sigma_{k}$ in (29) is fully Hadamard differentiable under some assumptions. For the precise definitions we refer to Cárcamo et al. (2020) and the references therein.
@@
Proof.
From (Cárcamo et al., 2020, Theorem 2.1), we have that $\sigma_{k}$ is Hadamard directionally differentiable and

### v3

@@ 4 Empirical results
In the same spirit as [12] or [23, Section 8.1], we emphasize the interest of a new homogeneity test (based on kernel distances), suitable for high-dimensional data and not suffering from the degradation of classical two-sample tests when the dimension increases. Additionally, we show the advantages of avoiding the choice of the parameters in kernel distances, via our SKD proposal. The general idea is to check the SKD methodology as an attempt to robustify the test statistic against bad choices of the kernel or its parameter(s). In this empirical study we compare the following methods:
@@
GKD: the kernel distance-based test of [23] with a data-driven choice of $\lambda$ in (2) as the median distance between points in the aggregate sample.
GKDSplit: test based on a Gaussian kernel distance where the estimation of the parameters is done by a splitting method; see [25]. The sample is divided into training and test subsamples to avoid data influence on the parameter selection.
GKDSplitOpt: test based on a kernel distance where the parameter estimation is detailed in [3, Section 3]. The data is divided into training and test sets. The target parameters are the coefficients of a convex combination of a finite family of kernels. The weights of the combination are selected to maximize the ratio between the empirical bias-corrected kernel distance and the standard deviation of the associated asymptotic distribution of the centered, normalized empirical distance.
ET: the energy test, a popular choice in this type of problems; see [41], [42].
In our view, the “data-splitting” proposals are based on natural ideas that deserve attention. Still, there are some open issues to clarify, especially regarding the optimal splitting of the sample and the asymptotic behaviour of the resulting data-driven tests. We hope that this study could encourage further research along these lines. As for the energy test ET ([41]), we have included it in the study because it is based in a successful statistical methodology, ultimately grounded on the underlying “distance covariance” association measure; in fact this method has become quite popular in high-dimensional two sample problems, via the energy package in R. Following a suggestion from one of the reviewers, we have considered as well a variant of this method, which is based on a different distance between the sample points. The standard statistic in ET is calculated in terms of the Euclidean distance, which can be seen, by a duality reasoning explained in [41], as a distance associated with the Brownian covariance. This suggests the possibility of broadening the choice of the distance to the whole range of the fractional Brownian motion with Hurst parameter $H\in(0,1)$; recall that $H={1}/{2}$ for the standard Brownian motion. In our case, the choice $H={3}/{4}$ led to results almost identical to those of $H={1}/{2}$. Perhaps if heavy-tailed distributions were involved, this new alternative could make a real difference.
@@
We include the models in [23], based on Gaussian distributions in high dimension with different means and diagonal covariance matrices. In addition, we also consider a new scenario with functional data corresponding to trajectories of Gaussian processes in $\operatorname{L}^{2}([0,1])$. We note that all the considered tests can be applied in the functional setting: GKD, SKD, GKDSplit and GKDSplitOpt are based on the aggregated matrix $\left(k_{\lambda}\left(Z_{i},Z_{j}\right)\right)_{i,j=1}^{n+m}$, where $Z_{l}=X_{l}$ for $l\in\{1,\ldots,n\}$ and $Z_{n+l}=Y_{l}$ for $l\in\{1,\ldots,m\}$. ET uses cross-distances between the data in the sample space.
More specifically, our simulation experiments are grouped in three blocks, respectively corresponding to different versions of homoscedasticity (Experiment 1) and heteroscedasticity (Experiment 2), plus a functional real data example.
@@
This regularization is common in harmonic analysis to approximate the Dirac delta in spaces of distributions via smooth functions, called mollifiers. It could be seen as an ad-hoc correction to improve the approximation of the maximum of the estimated kernel distance to the corresponding “true” population maximum. The smoothing process can be eliminated when sample sizes are sufficiently large.
@@
As shown in the literature, a data-driven choice of $\lambda$ seems to have a good practical behaviour. Specifically, in [23] (and in subsequent works), the value of $\lambda$ is the median distance between points in the aggregate sample. A theoretical consequence of this choice is that the asymptotic theory, derived in [23] under the assumption that $\lambda$ is fixed, does not longer apply to the data-driven case. Still, we include in our experiments, for comparison purposes, this data-driven choice as it is a common practice in the earlier literature. Since, to the best of our knowledge, the asymptotic distribution of the data-driven statistic is not known, we use a permutation test based on this statistic to obtain rejection regions rather than the other methods (Pearson curves, Gamma curves and bootstrap for U-statistics) explained in [23].
@@
It is worth mentioning that in both tests SKD and ET a permutation procedure has been used to approximate the corresponding distributions. This is the methodology used in the energy R-package for the ET test and we have followed here the same strategy for the SKD test. Let us recall that our theoretical results provide the asymptotic distribution of this test, as well as its consistency (see Theorems 2 and 3). However, the estimation of the quantiles of the limit distribution in (13) is far from trivial. As an additional complication, standard bootstrap approximation fails, as a consequence of the results in [22]. This is why the permutation method appears as a natural choice. Further, the validity of the permutation test comes from the permutation invariance of the test statistic (see [37]).
@@
Let us recall that the idea behind the SKD test is to dodge the parameter selection problem by considering “the whole parameter space”. Ideally, for the Gaussian kernel, an interval of the form $(0,\infty)$ could be considered in the SKD. As mentioned before, the extremes (00 and $\infty$) are not useful since the distance tends to zero when the parameter approaches to these values. Therefore, we use a parameter space of the form $\Lambda=[a,b]$, with $0<a<b<\infty$. It is important to note that here the values $a$ and $b$ cannot be properly considered as tuning parameters, since the test is not particularly sensitive to their choice, provided that the interval is large enough. As we have experimentally verified, $\Lambda=\left[10^{-4},0.1\right]$ is adequate to carry out the test. In practice, we employ a grid of $11$ points logarithmically separated between $10^{-4}$ and $0.1$. The goal is to approximate the value of the supremum by the maximum over finite subsets. The simulation outputs below are based on averages over $200$ replications. The permutation tests for GKD and SKD correspond to $B=5000$ permutations. As for ET, we use the function eqdist.etest of the R-package [42]. Sample sizes are $n=m=250$ in all experiments. The effect of increasing the dimension $d$ is checked in the rank $d\in\{205,405,603,803,1003,1203,1401,$ 1601,1801,2001}1601,1801,2001\}1601 , 1801 , 2001 }. In all cases, the significance level of the test is set at $\alpha=0.05$.
@@
We consider daily values of maximum temperatures registered at Barcelona airport (El Prat) from years 1944 to 2019. The data set consists of 76 vectors of dimension 365, each of which corresponds to a year in that time period. The daily observations have been treated as discretization points to include the problem within the framework of functional data, every year providing a function in the sample. Those observations corresponding to the 29th of February in leap years are omitted and missing observations are interpolated. These data are available at https://www.ncei.noaa.gov, the web page of the National Centers for Environmental Information.
Our purpose is to test the null hypothesis that the sample of temperatures from 1944 to 1981 comes from the same (functional) distribution to that of the period 1982-2019. The rejection of this null hypothesis could be interpreted as a hint of possible warming in the area. Indeed, we observe that, in absence of any significant climate change, one would expect that both samples are made of independent trajectories from the same underlying process.
@@
In the light of the results, we can conclude that, globally, the supremum kernel distance test (SKD) performs similarly to the GKD tests in the homoscedastic case, though the ET test appears to be the winner in this situation. In the heteroscedastic case, SKD obtains almost the best power results. A more complete study (including the derivation of the asymptotic distribution for the case of a data-driven selection of $\lambda$, the use of Pearson curves and/or modified bootstrap schemes, …) might be worthwhile in the future. On the other hand, according to [28], energy tests can be expressed in terms of kernel distances. This idea might deserve further attention as well, in order to incorporate these “equivalent” kernels to the SKD paradigm. In any case, it is clear that the present study does not allow us to conclude any obvious superiority (or inferiority) of none of the considered methods. In fact, the aim of our limited empirical study is to show that the SKD method can be implemented and it is competitive. This goal has been hopefully achieved. More definitive conclusions should be reached via subsequent empirical experiments and, especially, with the use of these tests by practitioners in the coming years. Software to run the SKD-based test will soon be available as an R-package called SKD2.
@@ 5 Final remarks
The whole paper relies on the strategy of using kernel-based methods by combining, via supremum, different kernels. We consider the classical two-sample problem focusing on functional and high-dimensional data, where the demand for applicable methods seems more obvious. Despite the large amount of relevant literature on this topic, we believe that there is still room for improvement, as those provided here, in the line of obtaining more general results with a different technology of proofs. In particular, the use of differentiation techniques plus empirical processes methods allows us to address the asymptotic behaviour of the test statistics under the null and alternative hypothesis, including the case of unbalanced samples; see Theorems 2 and 3. A key result in the proofs is Theorem 1 that establishes, under some conditions covering infinite-dimensional settings, the universal Donsker property for a union of unit balls in RKHS. This theorem extends previous similar results for finite-dimensional situations and could be potentially useful in other statistical procedures within the RKHS framework. The approach established in this paper can also be potentially useful to analyze the asymptotic behavior of data-driven estimators of the kernel parameters. However, this interesting problem is beyond the scope of this paper. Theorems 2 and 3 are meaningful from a conceptual point of view, even if the limit distributions are not particularly simple. The mere existence of such limit (non-degenerate) distributions is a primary guarantee that kernel-based statistics can be used to derive procedures achieving, at least asymptotically, a prescribed significance level under the null and providing consistency under the alternative. We deal with the complicated structure of the limit distribution under the null by using permutation tests. This is also the case for other popular methodologies such as the energy test, [41]. Different approximation techniques are also conceivable, including truncation in the limit expression in Theorems 2 plus estimation of the involved parameters. While this paper was under review, other related and interesting contributions have appeared; see [1] and [7].
@@ 6 Technical details
We need two auxiliary lemmata to prove Theorem 1. The first result corresponds to [38, Theorem 1.1].
@@
We also require Aronszajn’s inclusion theorem; see [2, Theorem I].
@@
The first part can be seen as a consequence of Lemma 1. First, by [6, Theorem 17], the functions in $\mathcal{H}_{k}$ are continuous. In particular, using [6, Corollary 3], we conclude that $\mathcal{H}_{k}$ is a separable Hilbert space. On the other hand, for $x\in\mathcal{X}$, by the reproducing property (6) of $k$ (twice) and Cauchy–Schwarz inequality, we have that
@@
Therefore, as $k$ is bounded on the diagonal, convergence in the RKHS norm entails uniform convergence. Further, from (19) we also see that the functions in $\mathcal{H}_{k}$ are bounded and hence $\mathcal{H}_{k}\subset\mathcal{C}_{b}(\mathcal{X})$. Now, we can apply Lemma 1 to $H=\mathcal{H}_{k}$ and $T=I$, the inclusion map given by $I(f)=f$. According to (19), this linear transformation is continuous. As $B_{H}=\mathcal{F}_{k}$, by Lemma 1, we thus conclude that $\mathcal{F}_{k}=T\left(\mathcal{F}_{k}\right)$ is universal Donsker.
The second part is a by-product of the first one together with Aronszajn’s inclusion theorem. According to Lemma 2, we have that $\|f\|_{\mathcal{H}_{k}}\leq\sqrt{c}\,\|f\|_{\mathcal{H}_{k,\lambda}}$, for all $f\in\mathcal{H}_{k,\lambda}$ and for all $\lambda\in\Lambda$. Therefore, $\mathcal{F}_{k,\Lambda}\subset\sqrt{c}\,\mathcal{F}_{k}$. Finally, from the first part of the theorem, the set $\sqrt{c}\,\mathcal{F}_{k}$ is universal Donsker as it is the unit ball of the RKHS generated by the kernel $ck$. Therefore, $\mathcal{F}_{k,\Lambda}$ is also universal Donsker (see [48, Theorem 2.10.1]) and the proof is complete. ∎
@@
Proof:.
To show part ((a)), we note that, from Theorem 1, $\mathcal{F}_{k,\lambda}$ is a $\operatorname{P}$-Donsker class and hence $\operatorname{P}$-pre-Gaussian. Hence, by [31, Theorem 3.7.28], for almost all $\omega$, the function $f\mapsto\mathbb{G}_{\operatorname{P}}(\omega)f$ ($f\in\mathcal{F}_{k,\lambda}$) is prelinear and can be uniquely extended to a linear map on $\operatorname{span}\left(\mathcal{F}_{k,\lambda}\right)=\mathcal{H}_{k,\lambda}$. Moreover, this extension is bounded and uniformly $d_{\operatorname{P}}$-continuous in $\mathcal{H}_{k,\lambda}$, where $d_{\operatorname{P}}$ is the intrinsic $\operatorname{L}^{2}(\operatorname{P})$ metric of the process. Finally, we observe that, thanks to (19),
@@
To prove part ((b)) we first note that, by ((a)), $\mathbb{G}_{\operatorname{P}}$ is a Gaussian process in the Hilbert space $\mathcal{H}_{k,\lambda}^{\ast}$. The covariance operator $\mathcal{K}_{\mathbb{G}_{\operatorname{P}}}$ of $\mathbb{G}_{\operatorname{P}}$ is self-adjoint and compact. By the Fernique’s theorem [9, p. 74], $\mathbb{G}_{\operatorname{P}}$ is Bochner square-integrable, $\mathcal{K}_{\mathbb{G}_{\operatorname{P}}}$ is a trace-class operator and
@@
where $\nu_{\mathbb{G}_{\operatorname{P}}}$ is the measure induced by the process $\mathbb{G}_{\operatorname{P}}$ in $\mathcal{H}_{k,\lambda}^{\ast}$. The proof of (23) can be found in [9, p. 48].
@@
From (24), we have that $Z_{j,\lambda}=\left\langle\mathbb{G}_{\operatorname{P}},\varphi_{j,\lambda}\right\rangle_{\mathcal{H}_{k,\lambda}^{\ast}}\sim\mathcal{N}\left(0,\beta_{j,\lambda}\right)$ ($j\in\mathbb{N}$) are jointly Gaussian and independent.
To finish this proof of (20), by [36, Theorem 6.1], it is enough to show absolute mean convergence, which is a necessary and sufficient condition. First, by orthogonality, we observe that for every $J\subset\mathbb{N}$ finite, we have that
@@
The proof of part ((a)) in Lemma 3 essentially follows from Theorem 1. However, part ((b)), where the series representation is obtained, must be discussed. Equation (20) shows the convergence of a series of functional random variables. This result looks like a standard Karhunen-Loève theorem, but some remarks should be done. The convergence of this series is on the dual space $\mathcal{H}_{k,\lambda}^{\ast}$, while Karhunen-Loève decomposition is stated classically on $\operatorname{L}^{2}$-type spaces. In fact, our decomposition in (20) can be seen as a particular case of the results in [5]. In [31, Theorem 2.6.10] a similar decomposition is shown where the coordinates are deterministic while the basis is random, which is not useful for our purposes.
@@
Note that $d_{k,\Lambda}$ is the metric induced by the supremum norm in $\ell^{\infty}\left(\mathcal{F}_{k,\Lambda}\right)$, hence $d_{k,\Lambda}$ is a continuous functional. From (26) and by the continuous mapping theorem (see, for instance [48, Theorem 1.9.5]), we obtain that
@@
The next goal is to prove Corollary 1 as preparation for the proof of Theorem 3. We need a differentiability result for the supremum similar to those obtained in [14]. Given a kernel $k$, we consider the mapping
@@
where $\mathcal{F}_{k,\Lambda}$ is the union of balls in (11). These differentiability results might have independent interest as it can be applied in other contexts by means of the (extended) functional Delta method; see the examples in [14].
The next corollary shows that $\sigma_{k}$ in (29) is fully Hadamard differentiable under some assumptions. For the precise definitions we refer to [14] and the references therein.
@@
Proof:.
From [14, Theorem 2.1], we have that $\sigma_{k}$ is Hadamard directionally differentiable and